OpenStudy (anonymous):

If 3 dice are thrown, what is the probability that the number appeared on one dice is equal to the sum of the numbers appeared on the other two dice?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@wolf1728

OpenStudy (anonymous):

If 3 dice are thrown, what is the probability that the number appeared on one dice is equal to the sum of the numbers appeared on the other two dice?

OpenStudy (anonymous):

2,(1,1) upto 6 (1,5) (3,3)(2,4),(4,2)(5,1)so total=15

OpenStudy (anonymous):

n(s)=6^3=216,so=15/216

ganeshie8 (ganeshie8):

this looks hard

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

the probab q

OpenStudy (ikram002p):

what is HCF ?

OpenStudy (anonymous):

probability question

ganeshie8 (ganeshie8):

I have edited the problem, check it now @ikram002p

OpenStudy (ikram002p):

If 3 dice are thrown, what is the probability that the number appeared on one dice is equal to the sum of the numbers appeared on the other two dice? *confused *

OpenStudy (anonymous):

if the number appeared in one dice is 2 then other two will be 2 i.e (sum of two dice is2) (1,1) if 3 (1,2) (2,1)

OpenStudy (ikram002p):

ok so we have 3 dice mmm 1,1 , 2 1,2 , 3 1,3 ,4 1,4,5 1,5,6 2,1 ,3 2,2,4 2,3,5 2,4,6 3,1,4 3,2,5 3,3,6 4,1,5 4,2,6 5,1,6

OpenStudy (ikram002p):

does order of the other two dice important ?

OpenStudy (anonymous):

no

OpenStudy (ikram002p):

ok so grouping (1,2,3,4,5) (1,2,3,4,5) =(1,1) (1,2) (1,3) (1,4) (1,5) (2,3) (2,4) (3,3) mmmm

OpenStudy (ikram002p):

P (of sum ) = p( of having sum 2 ) + p( of having sum 3) + p( of having sum 4) + P (of having sum 5) + p( of having sum 6)

OpenStudy (ikram002p):

p (of having sum 2) is 1/6 *1/6* 1/6 only when first dice =1 and second =1 and third = 2

OpenStudy (ikram002p):

p( of having sum 3 ) is 2/6 * 2/6 *1/6 (1,2,3) or (2,1,3)

OpenStudy (ikram002p):

@ganeshie8 mmm idk a bit confused

OpenStudy (anonymous):

Four-digit numbers are formed using the digits 0, 2, 3, 4, 5 without repetition. What is the probability that the numbers are divisible by 4?

OpenStudy (anonymous):

04,12,24,32,52 are last two numbers which is divisible by 4 so 5*3*2=30

OpenStudy (ikram002p):

what we are solving lol !!!

OpenStudy (ikram002p):

ic , so is it other Qn ?

OpenStudy (ikram002p):

ok i would rather to find whole space :P

OpenStudy (anonymous):

ya.i have lot of question to solve.its so much time for one question so atleast i can get answer to one question.its so boring if not getting answer

OpenStudy (ikram002p):

ok ok for first one i think finding the whole space which is 15/6^3

OpenStudy (ikram002p):

If 3 dice are thrown, what is the probability that the number appeared on one dice is equal to the sum of the numbers appeared on the other two dice? answer 15/6^3

OpenStudy (anonymous):

no.that's wrong .

OpenStudy (ikram002p):

:O why whats the right ansewer ?

OpenStudy (anonymous):

i don't kno

OpenStudy (ikram002p):

could it be like this ? (1/6)^3 + (2/6) (1/6)^2+ (3/6)(2/6)(1/6)

OpenStudy (ikram002p):

15/6^3 if we ignore order but 9/6^3 for order

OpenStudy (anonymous):

i know options 1/24,5/216,5/24,5/36,something like that

OpenStudy (ikram002p):

ohh so 9/6^3=1/24 mmm xD * confused*

OpenStudy (anonymous):

four digit number

ganeshie8 (ganeshie8):

correct answer is http://www.wolframalpha.com/input/?i=45%2F216

ganeshie8 (ganeshie8):

i have just listed down all the possibilities, not sure if there is a clever way.. .

OpenStudy (ikram002p):

mm for which Qn ?

ganeshie8 (ganeshie8):

for the original main question, @dg2 kindly limit to one question per post so that there won't be any confusions :)

OpenStudy (ikram002p):

huh yeah lol its confusing me so ganesh with listing posibilities u took it like this ? (2,1,1) (3,1,2) (3,2,1) (4,1,3) (4,3,1) (4,2,2 ) ??

ganeshie8 (ganeshie8):

``` 1 : 1 1 2 2 : 1 2 1 3 : 1 2 3 4 : 1 3 2 5 : 1 3 4 6 : 1 4 3 7 : 1 4 5 8 : 1 5 4 9 : 1 5 6 10 : 1 6 5 11 : 2 1 1 12 : 2 1 3 13 : 2 2 4 14 : 2 3 1 15 : 2 3 5 16 : 2 4 2 17 : 2 4 6 18 : 2 5 3 19 : 2 6 4 20 : 3 1 2 21 : 3 1 4 22 : 3 2 1 23 : 3 2 5 24 : 3 3 6 25 : 3 4 1 26 : 3 5 2 27 : 3 6 3 28 : 4 1 3 29 : 4 1 5 30 : 4 2 2 31 : 4 2 6 32 : 4 3 1 33 : 4 5 1 34 : 4 6 2 35 : 5 1 4 36 : 5 1 6 37 : 5 2 3 38 : 5 3 2 39 : 5 4 1 40 : 5 6 1 41 : 6 1 5 42 : 6 2 4 43 : 6 3 3 44 : 6 4 2 45 : 6 5 1 ```

OpenStudy (anonymous):

yes @ganeshie8

ganeshie8 (ganeshie8):

if that helps in reverse engineering... @ikram002p

ganeshie8 (ganeshie8):

@dg2 - i suggest you move on with your other problems, we will tag you once we figure out some solution :)

OpenStudy (anonymous):

well i believe the equation is asking about the sum of 2 numbers

OpenStudy (anonymous):

no the difference combined

OpenStudy (anonymous):

shall i close n post a new question

ganeshie8 (ganeshie8):

sure :)

ganeshie8 (ganeshie8):

Solve in in integers : 1<=(a,b,c)<=6 a = b+c or b = c+a or c = a+b

OpenStudy (ikram002p):

have nothing else to say @ganeshie8 lol the order thing confuse me i only thought first dice is the one which should be the sum so just multiply pre approch with 5 5 ( 1/6^3+2/6 1/6^2+ 3/6 2/6 1/6 )

ganeshie8 (ganeshie8):

thats working xD how did u get that ? care to explain it step by step

OpenStudy (ikram002p):

i know will possibilty without ordering is 1/6^3+2/6 1/6^2+ 3/6 2/6 1/6 so we have 5 orderd since only sum we can get 2,3,4,5,6

ganeshie8 (ganeshie8):

how this works : 1/6^3+2/6 1/6^2+ 3/6 2/6 1/6 ?

ganeshie8 (ganeshie8):

break it down please, my brain is malfunctioning today >.<

OpenStudy (larseighner):

It cannot be the case that more than one die is equal to the sum of the other two. So there is either 1 die that is the sum of the other two or 0 dice have that property. If there is 1 die, it must be the case that it is the one showing the greatest value. That value must be at least 1 more than the next largest die (or dice in case of a tie for second place).

OpenStudy (ikram002p):

look pre comments \(\uparrow\) p= p( of having sum 2 ) + p( of having sum 3) + p( of having sum 4) + P (of having sum 5) + p( of having sum 6) = 1/6^3+2/6 1/6^2+ 3/6 2/6 1/6 huh its something i tool in Abstract algebra about order grouping is we have odd set like (2,3,4,5,6) ( set of sum ) its of order 5 so directly in propabilty we can iterept it like this :- p( of having sum 2 ) = 1/6 * 1/6 * 1/6 p( of having sum 3) =1/6* 1/6*1/6 p( of having sum 4) =2/6 *1/6*1/6 p(of having sum 5) = 2/6*1/6*1/6 p(of having sum 6) =3/6*1/6*1/6

OpenStudy (larseighner):

So all three dice tying is out. The chance of getting a number on one die is 1. The chance of matching that number on the second die is 1/6 and of matching on the other die is 1/6. So the chance of hitting a triple is 1 x 1/6 x 1/6 = 1/36 so put 1/36 in the NOT category.

ganeshie8 (ganeshie8):

got that Lars, only one die equals the sum of other two. This follows directly from below : if a = b+c, then b = a-c, c = a-b

OpenStudy (anonymous):

ok how many possible combinations are there?

ganeshie8 (ganeshie8):

Oh that brilliant! so you want to work out the complement first is it

OpenStudy (anonymous):

yes

OpenStudy (ikram002p):

u can get sum of 2 form 1 compination sum of 3 from 1 compination sum of 4 from 2 compination sum of 5 from 2 compination sum of 6 from 3 compination

OpenStudy (larseighner):

I'm not liking this tak. How about the sum of the dice must be an even number?

OpenStudy (anonymous):

it says sum of two dices

OpenStudy (larseighner):

If the number on one die equals the sum on the other two dice, then the sum on all three dice must be even.

OpenStudy (larseighner):

If the greatest die is odd, then there must be exactly one other odd die.

OpenStudy (larseighner):

Case of greatest die is odd: Die is 5, exactly one of the other dice must be 3 or 1 Die is 3, exactly one of the other dice must be 1

OpenStudy (larseighner):

If the greatest die is even: both the other dice must be even or both must be odd.

ganeshie8 (ganeshie8):

beautiful !!

OpenStudy (larseighner):

Let's back up a little bit. I will make the bold claim that the chance that the sum of three dice summing to an odd number is .5. Can we prove that?

ganeshie8 (ganeshie8):

odd cases : 5 : (3, 2) , (1, 4) : 12 ways 3 : (1, 2) : 6 ways even cases : 6 : (1, 5), (3, 3), (2, 4), : 15 ways 4 : (1, 3), (2, 2) : 9 ways 2 : (1, 1) : 3 ways ------------------------------ 45 ways this looks to be the shortest method possible xD

OpenStudy (ikram002p):

wow !

ganeshie8 (ganeshie8):

we can prove that claim easily - the sum of 3 numbers is either odd or even, by symmetry we can conclude that the chances of getting sum odd is 1/2

OpenStudy (ikram002p):

this is how it works in odd order grouping xD

OpenStudy (anonymous):

i see

OpenStudy (larseighner):

Okay, I see the proof is more or less trivial. The chances of the total (of three) being odd is .5 which is OUT Now we figured the chance of all being tied at 1/36 but half of those are odd totals already excluded. So three tied and even = 1/72 which is OUT.

OpenStudy (larseighner):

Humm running out of gas here. I'll look back and see how things are going the other way.

OpenStudy (larseighner):

216 is the total cases. I eliminated half of them leaving 108. Then I eliminated another 1/72 of them, leaving 105. @ganeshie8 accounts for 45 ways of winning. So I am looking for a genius way of eliminating 60 cases.

ganeshie8 (ganeshie8):

108 - odd 3 - even and all dice equal ---------------- 111 ways are eliminated so far, leaving 105 ways

OpenStudy (larseighner):

Random hypotheses: If exactly two dice are tied -> neither can be the greatest die and the greatest must be even.

ganeshie8 (ganeshie8):

yup! two dice can be tied in 6 ways

ganeshie8 (ganeshie8):

oh wait, the converse is not true

OpenStudy (larseighner):

It cannot be true that the top die is 6 and only one of the others is odd. It cannot be true that the top die is 5 and both of the others is even. It cannot be true that the top die is 4 and only one of the others is odd. It cannot be true that the top die is 3 and both of the others is even. It cannot be true that the top die is 2 and only one of the others is odd. The case that the top die is 1 was eliminated by excluding odd totals of all three dice.

OpenStudy (larseighner):

Hmm...bored to death, and the drunk woke up and is talking to space aliens.